OpenStudy (anonymous):

Solve by any appropriate method (substitution or elimination) 12x+y=465 12y+x=146 @Loser66 @e.mccormick

OpenStudy (loser66):

I prefer substitute. try, please

OpenStudy (anonymous):

I too..

OpenStudy (loser66):

help her, please @waterineyes

OpenStudy (anonymous):

ok. I understand this one a little better, but the only difference with this one than the ones I've gotten use to is the position of y and x in each equation.

OpenStudy (anonymous):

Here find y from first equation: you will get : \(12x + y = 465\) Now subtract 12x from both the sides: \(y = 465 - 12x\)..

OpenStudy (anonymous):

Now this value you will use in second equation in place of y, and solve that you will get x..

OpenStudy (anonymous):

so the second equation will look like; 12(465-12x)+x=146?

OpenStudy (anonymous):

5580-144x+x=146 5580-143x=146 143x=38.219\[\approx38.22\]

OpenStudy (anonymous):

Wait.. First two steps are right, what you have done in third??

OpenStudy (anonymous):

\(5580 - 143x = 146\) Here subtract 146 from both sides first and tell me what will you get?

OpenStudy (anonymous):

the 143 isn't supposed to be there. It's supposed to be x=38.219 but round it to 38.22

OpenStudy (anonymous):

143x is right, but firstly subtract 146 from both the sides..

OpenStudy (anonymous):

but that wouldn't leave me anything to find x

OpenStudy (anonymous):

See the step goes like this: \[5580 - 143x = 146\] Subtract here 146: \[5580 - 143x - 146 = 146 - 146\] \[5434 - 143x = 0\] Now add 143x both the sides: \[5434 - 143x + 143x = 0 + 143x\] \[5434 = 143x \; \; Or \; \; 143x = 5434\] Now divide it by 143 and fine x.

OpenStudy (anonymous):

oh. ok. That gives me 38

OpenStudy (anonymous):

Yep. Now find y by using \(y = 465 - 12x\)

OpenStudy (anonymous):

You will get a number starting from "N" when said in English language.. And you are done I think.. :)

OpenStudy (anonymous):

I'm so confused now I thought I understood this to a decent point but now I'm back to square one and don't understand it at all

OpenStudy (anonymous):

Where are you not getting?? You have got x, no??

OpenStudy (anonymous):

it's going to the second equation now to get y.

OpenStudy (anonymous):

See by plugging in value of x in any equation either first or second, you can find y..

OpenStudy (anonymous):

In any one of the equation, put x = 38, and find y..

OpenStudy (anonymous):

\[y=465-12*38\]\[y=465-456\]\[y=9\]\[x=38 y=9\] @waterineyes

OpenStudy (anonymous):

Yep.. Good..

OpenStudy (anonymous):

ok. Thank you

OpenStudy (anonymous):

The thing that matters is not this that you got right answer but have you got all the steps, how we do it.. Confirm it.. You are welcome dear..

OpenStudy (anonymous):

I do understand. I had to go back and relook at it. I missed writing down a couple things and that is what was frustrating me as far as not understanding. Thank you

OpenStudy (anonymous):

Okay, keep it up.. If still in problem, then let me know..

OpenStudy (anonymous):

I got that one. I have another one that deals with graphing and I understand how to do slope but I need a refresher on it